OpenStudy (yasmin):

(x2 + 3x + 4)(2x2 + 4x + 7)

OpenStudy (jdoe0001):

hmmm oooooooooook

OpenStudy (asc.bchs):

factor them out? set to zero and solve for x? multiply them? you need to be specific with what to do.

OpenStudy (jdoe0001):

if you're asking on how to multiply them well, take each term on one, and multiply each term by the other then combine their results say for example \(\bf \begin{array}{llll} x^2 + 3x + 4 \\ \quad \\ \qquad \times \quad 2x^2 \\\hline\\ \square ?+\square ?+\square ? \end{array}\quad \begin{array}{llll} x^2 + 3x + 4 \\ \quad \\ \qquad \times \quad 4x \\\hline\\ \square ?+\square ?+\square ? \end{array}\quad \begin{array}{llll} x^2 + 3x + 4 \\ \quad \\ \qquad \times \quad 7 \\\hline\\ \square ?+\square ?+\square ? \end{array}\) then combine their results

OpenStudy (sunnnystrong):

\[(x^2+3x+4)(2x^2+4x+7)=2x^4+10x^3+27x^2+37x+28\] =

OpenStudy (sunnnystrong):

@yasmin .. yes what is the goal of this problem XD

OpenStudy (mathmale):

Yasmin, the goal of the problem is to complete the indicated multiplication.

OpenStudy (mathmale):

Multiply (x^2 + 3x + 4) by (2x^2 + 4x + 7), or vice versa. The first of the 3 terms in the first multiplicand is x^2. Multiply each of the 3 terms in the 2nd multiplicand by x^2. Next, by 3x. Next, by 4. Add all the resulting products together, combining like terms.